OpenStudy (anonymous):

Solve. 2x^2 + 12x = 10 –3 ± √14 –3 ± 2√2 –3 ± √19 –3 ± √29

OpenStudy (nurali):

i think A http://www.wolframalpha.com/input/?i=2x%5E2+%2B+12x+%3D+10

OpenStudy (anonymous):

set it up as ax^2+bx+c=0 to do that you would subtract 10 and get 2x^2+12x-10=0. if i were you i would divide by 2 to get the leading coefficient as 1 x^2+6x-5=0 from there factor.